Kingston anticipates total sales for June and July of $340,000 and $378,000, respectively. Cash sales are normally 60% of total sales. Of the credit sales, 25% are collected in the same month as the sale, 65% are collected during the first month after the sale, and the remaining 10% are collected in the second month after the sale. Determine the amount of accounts receivable reported on the company’s budgeted balance sheet as of July 31.
Answer #1

Solution:

Accounts receievable as of July 31 for June sales = June credit sales*10% = $340,000*40%*10% = $13,600

Accounts receievable as of July 31 for July sales = July credit sales*75% = $378,000*40%*75% = $113,400

Total accounts receivable as of July 31 = 13600+ 113400 = $127,000